Today, Mr. D is 21 years younger than his neighbor. 3 years ago, the neighbor was twice as old as Mr. D was back then. Find the neighbor's current age. Please provide a chart or let statement.
Answer by ankor@dixie-net.com(22740)

Today, Mr. D is 21 years younger than his neighbor. 3 years ago, the neighbor was twice as old as Mr. D was back then. Find the neighbor's current age.
:
Let d = D's present age
Let n = neighbor's present age
:
" Mr. D is 21 years younger than his neighbor."
d = n - 21
:
3 years ago, the neighbor was twice as old as Mr. D was back then.
n - 3 = 2(d - 3)
n - 3 = 2d - 6
n = 2d - 6 + 3
n = 2d - 3
Replace d with (n-21)
n = 2(n-21) - 3
n = 2n - 42 - 3
n - 2n = -45
-n = - 45
n = 45 yrs neighbor's present age
;
:
See if that flies in the statement (d = 45 - 21 = 24);
3 years ago, the neighbor was twice as old as Mr. D was back then.
45 - 3 = 2(24 - 3)
42 = 2(21)